OpenStudy (anonymous):

find the slope of the line through each pair of points (Please show your work!) (12, -18), (11,12)

myininaya (myininaya):

\[\frac{12-(-18)}{11-12}=\frac{20}{-1}=-20\] this is the one i said was suppse to be -20 do you not get it?

OpenStudy (anonymous):

i did i just need all work shown :)